Need help to solve these questions with explanation. Pllease solve all questions please

12. What is of formations provided? nolar enthalpy change AHnu for the following reactions using the standard molar enthalpies sich() + 2 11,0 (l) ? SiO2 (s) + 4 HCl (g) AH (SiCIA) -640.1 kJ/mol AH (SiO2)-910.9 kJ/mol AH (H2O) -285.8 kJ/mol AH (HCI)-92.3 kJ/mol a. -2491.8 kJ/mol b. -1929.1 kJ/mol c. -77.3 kJ/mol d. -68.4 kJ/mol e. None of these 13. Given the following thermochemical equations and their corresponding enthalpies: 2 C H2 (g) +5 02(g) 4 CO2 (g) +2 H,o() CsHs (t)+10 02 (g)-8 CO2 (8)+4 H2O () AH -2601 kJ ??_-4393 kJ What is the enthalpy change for the following reaction? 4 C2H2 (g)-C8Hs () AH-? a. -1792 kJ b. -809 kJ C. -6994 kJ d. -9595 k.J e. none of these 14. Evaluate Af for the following reaction from the given bond energies. 2HBr(g) ->H2(g)+ Br2g) AHH-H-436 kJ/mol, Afr-Br-193 kJ/mol, AHH-Br-366 kJ/mol a. -103 k.J b. -143 kJ c. +103 kJ d. +142 kJ e. 259 kJ
Answer #1

1) SiCl4 + 2H2O -------> SiO2 + 4HCl

\DeltaHf(SiCl4) = -640.1kJ/mol \Delta Hf(H2O) = -258.8 kJ/mol \Delta Hf(SiO2) = -910.9kJ/mol   \Delta Hf(HCl) = -92.3kJ/mol

\DeltaHrxn = \sum n \Delta Hf(products) - \sum n \Delta Hf(reactants)

\sum n \Delta Hf(products) = -910.9+ 4(-92.3) = -1280.1 kJ/mol

\sumn \Delta Hf(reactants) = -640.1+ 2(-285.8) = -1211.7 kJ/mol

\DeltaHrxn = -1280.1 - (-1211.7) = -68.4 kJ/mol

2) 2C2H2 + 5O2 -----------> 4CO2 + 2H2O   \DeltaH = -2601 kJ ------- MULTIPLY BY 2

C8H8 + 10O2 -----------> 8CO2 + 4H2O   \DeltaH = - 4393 kJ ------ REVERSE THE REACTION

2( 2C2H2 + 5O2 -----------> 4CO2 + 2H2O)   \DeltaH = 2(-2601) kJ

4C2H2 + 10O2 -----------> 8CO2 + 4H2O    \DeltaH = -5202 kJ ---- 1

8CO2 + 4H2O -----------> C8H8 + 10O2\DeltaH = +4393 kJ ----- 2

Adding equation 1 and 2 gives the equation 3

   4C2H2(g) ------> C8H8(l) --------------- 3  \DeltaH = ? kJ

\DeltaH = \Delta Heq1 + \Delta Heq2 = -5202+4393 = -809 kJ

3) 2HBr -----> H2 + Br2

\DeltaHf(H-H) = 436 kJ/mol \Delta Hf(Br-Br) = 193 kJ/mol \Delta Hf(H-Br) = 366 kJ/mol

\DeltaHrxn = \sum n \Delta Hf(products) - \sum n \Delta Hf(reactants)

\sum n \Delta Hf(products) = 436 + 193  = 629 kJ/mol

\sumn \Delta Hf(reactants) = 2(366) = 732 kJ/mol

\DeltaHrxn = 732 - 629 = +103 kJ/mol